# A Priori Assessment of Rotational Invariance in Multiscale CNN-Based Subgrid-Scale Model for Wall-Bounded Turbulent Flows

### v1

@@ III Verification of material objectivity
For the non-rotated input, all models successfully reproduce both $\tau_{11}$ and $\tau_{12}$, indicating that the data-driven SGS models can capture the correct range of the subgrid stress magnitudes, despite an underprediction of both tails of the $\tau_{12}$ distribution. However, consistent with the wall-normal distributions in Figs. 7 and 9, significant differences emerge when the input fields are rotated. The MSC model exhibits noticeable deviations from the fDNS data, particularly for $\tau_{11}$, where the predicted distribution shifts toward lower values and loses the characteristic positive tail observed in the fDNS results. As previously discussed, the degradation in $\tau_{11}$ prediction can be attributed to its strong dependence on the input orientation. When the coordinate frame is rotated, the streamwise direction is no longer aligned with the trained reference frame, leading to a mismatch between the learned mapping of the input features and the corresponding stress component. For $\tau_{12}$, a noticeable discrepancy also appears in the MSC model, which fails to reproduce the fDNS distribution. In addition, the MSC2 model shows a tendency to underpredict the number of occurrences of $\tau_{11}$ and $\tau_{12}$, suggesting a suppression of the SGS effects. This may lead to reduced turbulence interactions and weakened energy transfer between the resolved and subgrid-scales, which likely results in an underestimation of SGS dissipation. A prominent improvement is shown by the MSC3 model, which reproduces both $\tau_{11}$ and $\tau_{12}$ more accurately for the rotated input. However, for $\tau_{12}$, it shows discrepancies at both tails, where the negative tail is strongly overpredicted at $90^{\circ}$ rotation. This may be attributed to limitations in the STN algorithm, particularly its reduced ability to generalize the rotated input, possibly owing to limited exposure to the rotated input during training. Such an overprediction of the negative shear stress may distort the energy transfer, particularly near the wall, where it may lead to a reduction in the SGS dissipation. From these findings, it can be concluded that although the MSC2 model can represent rotational invariance reasonably well through architectural constraints, achieving a higher accuracy requires additional augmentation, such as the inclusion of the STN algorithm, as depicted in the MSC3 model.
@@
Figure 11: Wall-normal distributions of SGS transport $\varepsilon_{\mathrm{sgs}}$ with input rotated by (a) $0^{\circ}$, (b) $90^{\circ}$, and (c) $270^{\circ}$.
@@
To further evaluate the data-driven SGS models quantitatively, statistical comparisons are presented to analyzed the underlying physical behavior captured by each model. Following the same procedure described in the preceding section, unseen data with rotated inputs of $90^{\circ}$ and $270^{\circ}$ are employed. Fig. 11 illustrates the SGS dissipation, defined under the local equilibrium assumption as $\varepsilon_{\mathrm{sgs}}=-\tau_{ij}\bar{D}_{ij}$. For a non-rotated input, the MSC and MSC2 models exhibit comparable accuracy, except near the peak region. In contrast, the MSC3 model shows a closer agreement with the fDNS data across the wall-normal direction. Because peak regions are often associated with intense turbulent dynamics and high spatial variability, their accurate prediction is challenging. It is likely that the integration of artificial modifications, particularly the inclusion of the STN algorithm in the MSC3 model, enhances the accuracy of the model in this region. During training, this model was exposed to multiple rotated versions of the input, allowing the STN to infer and correct for orientation differences. This mechanism functions similarly to data augmentation, as discussed by Brener et al. brener2024, but is more efficient Jaderberg, as it does not increase the amount of training data and is integrated directly into the backpropagation process within the data-driven SGS framework. These augmentations enable the model to better capture the turbulent interactions and improve its ability to predict the flow characteristics, even for non-rotated inputs.
@@
Figure 12 presents the SGS backscatter, defined as $\varepsilon_{b}=-\frac{1}{2}\langle\varepsilon_{\mathrm{sgs}}-|\varepsilon_{\mathrm{sgs}}|\rangle$, which represents the reverse transfer of energy from the subgrid scales to resolved scales. Here, the MSC2 and MSC3 models show considerably better alignment with the fDNS data when compared with that of the MSC model in the case of a non-rotated input. The MSC model slightly overpredicts the backscatter, which is likely owing to the spatial distortion introduced by BN. The BN may distort the spatial interaction, which then influences the local interaction on the backward cascade. Despite these slighly overestimated results, all models considered here could produce comparable levels of backscatter, suggesting that modeling the multiscale interactions in the data-driven SGS model based on the MSC model effectively captures the backscatter, as noted by Jalaali &\& Okabayashi jalaali2025.
@@
Next, for the SGS transport term in Fig. 13, defined as $T_{\mathrm{sgs}}=\partial(\tau_{ij}\bar{u}_{i})/\partial x_{j}$, all models perform comparably well for the non-rotated input. The SGS transport term describes the energy transfer within LES systems volker2002, and the accurate prediction of this quantity is crucial. These results show that all models exhibit almost no significant differences in the prediction performance. In the near-wall region ($y^{+}<20$), all models successfully reproduce the negative peak, followed by a nearly zero value away from the wall. These results indicate that the data-driven SGS models preserve the correct near-wall behavior observed in the fDNS data.
@@ IV Conclusion
To ensure physical consistency, a data-driven SGS model must satisfy the invariance principles underlying turbulence modeling or the principle of material objectivity, as pointed out by Spalart spalartphilosophies2015. Therefore, in this study, we proposed a novel data-driven SGS model for LES, with a focus on enforcing rotational invariance for a wall-bounded flow. Building upon the multiscale CNN-based SGS framework proposed by Jalaali & Okabayashi jalaali2025 (MSC model), we introduced two novel data-driven SGS models, denoted as the MSC2 and MSC3 models in this paper. The MSC2 model enforced material objectivity by excluding the bias terms and BN layers, whereas the MSC3 model further incorporated an STN algorithmJaderberg to learn the spatial alignment under the rotated input.
@@ Acknowledgements.
This work was partly supported by the Research Proposal-based Use of the Project for Nurturing Student Competing with the World at the Large-Scale Computer System \-- D3 Center, The University of Osaka. This study was financially supported by JSPS KAKENHI grant No. JP25K07584 and a grant from Beasiswa Pendidikan Indonesia (BPI)/ LPDP (the Indonesian Endowment Fund for Education, Ministry of Finance of the Republic of Indonesia).
@@
Parameter MSCjalaali2025 MSCa MSCb MSCc MSC2 MSC3
@@
Appendix A Assessment of the effect of bias, BN, and STN algorithm on the objectivity in a priori test
In this section, we report the effects of excluding the bias and batch normalization (BN), as well as introducing the spatial transformer network (STN) algorithmJaderberg in a data-driven SGS model, as determined through a priori tests. The exclusion of bias and BN has been shown to improve the objectivity of data-driven SGS models, yet their respective contributions to this improvement remain unclear. Therefore, a dedicated evaluation of these architectural components is essential to identify the architectural components of the DNN that most significantly influence the ability of the data-driven SGS model to accurately predict $\tau_{ij}$ and infer the objectivity. Here, all models were trained using the same dataset (i.e., the total number of training datasets was kept identical across all models, with MSCc and MSC3 employing rotated data, as described in Section II) and training procedure, ensuring a fair and consistent comparison. The model configurations used in this study are summarized in Table 2. To evaluate both the predictive performance and rotational invariance of the models, we examined the wall-normal distribution of the shear stress component $\tau_{12}$.
@@ Appendix A Assessment of the effect of bias, BN, and STN algorithm on the objectivity in a priori test
For the MSCa model, the exclusion of bias does not produce any significant effect on the non-rotated input, as it preserves its predictive accuracy. However, it exhibits noticeable deviations when tested under the rotated input. This result is expected, as the presence of BN alters the structure of the spatial correlations. By normalizing the output of the CNN layers across mini-batches, BN implicitly partitions the training data. We argue that this partitioning breaks the spatial coherence and disturbs the underlying vortex interaction, which can lead to unphysical representations. Turbulent flows are closely related to the local interaction and spatial coherencedavidson2015, and the exclusion of BN is critical when constructing data-driven SGS models. In addition, in the MSC model, alteration in the flow representation owing to BN appears to be corrected by bias. This result is consistent with the observations from the MSCb model, where the exclusion of BN enables it to better predict the rotated input. Because the input variable $\bar{D}_{ij}$ is inherently objective, the MSCb model benefits from this property and is able to develop a more rotation-invariant representation. However, the inclusion of bias in the MSCb model reintroduces an offset into the network, which leads to a degradation in accuracy and failure to satisfy the material objectivity requirement when tested under rotated input. The behaviors of the MSC2 and MSC3 models are discussed in the main section, where their performances under both rotated and non-rotated conditions have been analyzed in detail. Therefore, further discussion of the MSC2 and MSC3 models is omitted here.

### v2

@@ III Verification of material objectivity
For the non-rotated input, all models successfully reproduce both $\tau_{11}$ and $\tau_{12}$, indicating that the data-driven SGS models can capture the correct range of the subgrid stress magnitudes, despite an underprediction of both tails of the $\tau_{12}$ distribution. However, consistent with the wall-normal distributions in Figs. 7 and 9, significant differences emerge when the input fields are rotated. The MSC model exhibits noticeable deviations from the fDNS data, particularly for $\tau_{11}$, where the predicted distribution shifts toward lower values and loses the characteristic positive tail observed in the fDNS results. As previously discussed, the degradation in $\tau_{11}$ prediction can be attributed to its strong dependence on the input orientation. When the coordinate frame is rotated, the streamwise direction is no longer aligned with the trained reference frame, leading to a mismatch between the learned mapping of the input features and the corresponding stress component. For $\tau_{12}$, a noticeable discrepancy also appears in the MSC model, which fails to reproduce the fDNS distribution. In addition, the MSC2 model shows a tendency to underpredict the number of occurrences of $\tau_{11}$ and $\tau_{12}$, suggesting a suppression of the SGS effects. This may lead to reduced turbulence interactions and weakened energy transfer between the resolved and subgrid-scales, which likely results in an underestimation of SGS dissipation. A prominent improvement is shown by the MSC3 model, which reproduces both $\tau_{11}$ and $\tau_{12}$ more accurately for the rotated input. However, for $\tau_{12}$, it shows discrepancies at both tails, where the negative tail is strongly overpredicted at $90^{\circ}$ rotation. This may be attributed to limitations in the STN algorithm, particularly its reduced ability to generalize the rotated input, possibly owing to limited exposure to the rotated input during training. Such an overprediction of the negative shear stress may distort the energy transfer, particularly near the wall, where it may lead to a reduction in the SGS dissipation. From these findings, it can be concluded that although the MSC2 model can represent rotational invariance reasonably well through architectural constraints. However, achieving a higher accuracy requires additional augmentation, such as the inclusion of the STN algorithm, as depicted in the MSC3 model.
@@
Figure 11: Wall-normal distributions of SGS dissipation $\varepsilon_{\mathrm{sgs}}$ with input rotated by (a) $0^{\circ}$, (b) $90^{\circ}$, and (c) $270^{\circ}$.
@@
To further evaluate the data-driven SGS models quantitatively, statistical comparisons are presented to analyzed the underlying physical behavior captured by each model. Following the same procedure described in the preceding section, unseen data with rotated inputs of $90^{\circ}$ and $270^{\circ}$ are employed. Fig. 11 illustrates the SGS dissipation, defined under the local equilibrium assumption as $\varepsilon_{\mathrm{sgs}}=-\tau_{ij}\bar{D}_{ij}$. For a non-rotated input, the MSC and MSC2 models exhibit comparable accuracy, except near the peak region. In contrast, the MSC3 model shows a closer agreement with the fDNS data across the wall-normal direction. Because peak regions are often associated with intense turbulent dynamics and high spatial variability, their accurate prediction is challenging. It is likely that the integration of artificial modifications, particularly the inclusion of the STN algorithm in the MSC3 model, enhances the accuracy of the model in this region. During training, this model was exposed to multiple rotated versions of the input, allowing the STN to infer and correct for orientation differences. This mechanism functions similarly to data augmentation, as discussed by Brener et al. [7], but is more efficient [20], as it does not increase the amount of training data and is integrated directly into the backpropagation process within the data-driven SGS framework. These augmentations enable the model to better capture the turbulent interactions and improve its ability to predict the flow characteristics, even for non-rotated inputs.
@@
Figure 12 presents the SGS backscatter, defined as $\varepsilon_{b}=-\frac{1}{2}\langle\varepsilon_{\mathrm{sgs}}-|\varepsilon_{\mathrm{sgs}}|\rangle$, which represents the reverse transfer of energy from the subgrid scales to resolved scales. Here, the MSC2 and MSC3 models show considerably better alignment with the fDNS data when compared with that of the MSC model in the case of a non-rotated input. The MSC model slightly overpredicts the backscatter, which is likely owing to the spatial distortion introduced by BN. The BN may distort the spatial interaction, which then influences the local interaction on the backward cascade. Despite these slighly overestimated results, all models considered here could produce comparable levels of backscatter, suggesting that modeling the multiscale interactions in the data-driven SGS model based on the MSC model effectively captures the backscatter, as noted by Jalaali &\& Okabayashi [21].
@@
Next, for the SGS transport term in Fig. 13, defined as $T_{\mathrm{sgs}}=\partial(\tau_{ij}\bar{u}_{i})/\partial x_{j}$, all models perform comparably well for the non-rotated input. The SGS transport term describes the energy transfer within LES systems [43], and the accurate prediction of this quantity is crucial. These results show that all models exhibit almost no significant differences in the prediction performance. In the near-wall region ($y^{+}<20$), all models successfully reproduce the negative peak, followed by a nearly zero value away from the wall. These results indicate that the data-driven SGS models preserve the correct near-wall behavior observed in the fDNS data.
@@ IV Conclusion
To ensure physical consistency, a data-driven SGS model must satisfy the invariance principles underlying turbulence modeling or the principle of material objectivity, as pointed out by Spalart [41]. Therefore, in this study, we proposed a novel data-driven SGS model for LES, with a focus on enforcing rotational invariance for a wall-bounded flow. Building upon the multiscale CNN-based SGS framework proposed by Jalaali & Okabayashi [21] (MSC model), we introduced two novel data-driven SGS models, denoted as the MSC2 and MSC3 models in this paper. The MSC2 model enforced material objectivity by excluding the bias terms and BN layers, whereas the MSC3 model further incorporated an STN algorithm[20] to learn the spatial alignment under the rotated input.
@@ Acknowledgements.
This work was partly supported by the Research Proposal-based Use of the Project for Nurturing Student Competing with the World at the Large-Scale Computer System \-- D3 Center, the University of Osaka. This study was financially supported by JSPS KAKENHI grant No. JP25K07584 and Indonesian Education Scholarship Program (BPI), administered by the Center for Higher Education Funding and Assessment (PPAPT), Ministry of Higher Education, Science, and Technology of the Republic of Indonesia and the Indonesia Endowment Fund for Education (LPDP), Ministry of Finance of the Republic of Indonesia.
@@
Parameter MSC[21] MSCa MSCb MSCc MSC2 MSC3
@@
Appendix A Assessment of the effect of bias, BN, and STN algorithm on the objectivity in a priori test.
In this section, we report the effects of excluding the bias and batch normalization (BN), as well as introducing the spatial transformer network (STN) algorithm[20] in a data-driven SGS model, as determined through a priori tests. The exclusion of bias and BN has been shown to improve the objectivity of data-driven SGS models, yet their respective contributions to this improvement remain unclear. Therefore, a dedicated evaluation of these architectural components is essential to identify the architectural components of the DNN that most significantly influence the ability of the data-driven SGS model to accurately predict $\tau_{ij}$ and infer the objectivity. Here, all models were trained using the same dataset (i.e., the total number of training datasets was kept identical across all models, with MSCc and MSC3 employing rotated data, as described in Section II) and training procedure, ensuring a fair and consistent comparison. The model configurations used in this study are summarized in Table 2. To evaluate both the predictive performance and rotational invariance of the models, we examined the wall-normal distribution of the shear stress component $\tau_{12}$.
@@
For the MSCa model, the exclusion of bias does not produce any significant effect on the non-rotated input, as it preserves its predictive accuracy. However, it exhibits noticeable deviations when tested under the rotated input. This result is expected, as the presence of BN alters the structure of the spatial correlations. By normalizing the output of the CNN layers across mini-batches, BN implicitly partitions the training data. We argue that this partitioning breaks the spatial coherence and disturbs the underlying vortex interaction, which can lead to unphysical representations. Turbulent flows are closely related to the local interaction and spatial coherence[12], and the exclusion of BN is critical when constructing data-driven SGS models. In addition, in the MSC model, alteration in the flow representation owing to BN appears to be corrected by bias. This result is consistent with the observations from the MSCb model, where the exclusion of BN enables it to better predict the rotated input. Because the input variable $\bar{D}_{ij}$ is inherently objective, the MSCb model benefits from this property and is able to develop a more rotation-invariant representation. However, the inclusion of bias in the MSCb model reintroduces an offset into the network, which leads to a degradation in accuracy and failure to satisfy the material objectivity requirement when tested under rotated input. The behaviors of the MSC2 and MSC3 models are discussed in the main section, where their performances under both rotated and non-rotated conditions have been analyzed in detail. Therefore, further discussion of the MSC2 and MSC3 models is omitted here.